The spot rate between the Japanese yen and the U.S. dollar is ¥106.57/$, while the one-year forward rate is ¥105.73/$. The one-year risk-free rate in the U.S. is 2.39 percent. If interest rate parity exists, what is the one-year risk-free rate in Japan?
Answer #1

As Per Interest Rate Parity Theory

\large Forward Rate = Spot Rate * \frac{1+Rf(Japan)}{1+Rf(US)}

\large 105.73 = 106.57 * \frac{1+Rf(Japan)}{1+0.0239}

\large 105.73 = 104.08 * (1+Rf(Japan) )

1+Rf(Japan) = 105.73/104.08

1+Rf(Japan) = 1.0159

Rf (Japan) = 1.0159-1 = 0.0159 OR 1.59%

Risk Free Rate is japan should be 1.59% as per Interest Rate Parity Theory

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.